# DiMA: An LLM-Powered Ride-Hailing Assistant at DiDi

### v1

@@ Appendix A Appendix
A.1. Simulation Environment Construction
A.1.1. Dialog Content Generation
A.1.2. GPT-based User Role-playing
With the generated dialog content using collected online data, we can further prompt GPT-4o to play the role of a real-world user following the dialog content. Specifically, we first provide a task profile in the prompt to guide GPT-4o in understanding the task, and then we provide the dialog content and prompt it to generate the first sentence to begin a dialog. In such a multi-dialog scenario, the response of the assistant will also be iteratively fed into the prompt to help GPT-4o understand if the ride-hailing intent has been completed. We provide detailed prompt in Figure LABEL:Dialog_Player to facilitate the understanding of GPT-based user role-playing.
A.1.3. Instruction Dataset Statistics
Figure 7. The prompt template of GPT-based user role-playing.
@@ A.2. Prompt Template of DiMA
This section presents the prompt template of every module in DiMA to facilitate a better understanding.
A.2.1. Spatiotemporal-aware Tool-augmented Order Planning
As mentioned in the methodology, this module conduct spatial mapping, temporal reasoning and progressive order planning. The entire process is achieved through function call. In this section, we provide the detailed prompt template in Figure 8. As can be seen in the demonstration in the prompt , when the user just input a address Hongyuan, DiMA will automatically generate a set of function call, which will be sequentially executed to achieve order planning.
Figure 8. Prompt template of spatiotemporal-aware tool-augmented order planning.
Figure 9. Prompt template of dialog replier selection mechanism.
A.2.2. Multi-type Dialog Replier Selection
After order planning, the dialog policy will be generated to allocate different dialog replier for response generation. Specifically, we fully utilize all current information for allocating, including current user query, history dialog, current time, and intermediate executed system results. We provide prompt template in Figure 9. As can be seen, we first provide a well-illustrative description of this task, and then we explain the input respectively, including history dialog, current user query, executing results, current time and order status. Then, we provide description of each dialog replier scenario to help it understand how to perform response replier allocation. Before controlling output format, we add demonstration example in the prompt.
Figure 10. Prompt template of specialized dialog replier.
A.2.3. Specialized Dialog Relpier
In specialized response generation pipeline, we will follow a formulaic response style to generate the assistant’s response. The prompt template is provided in Figure 10. As can be seen, we provide four illustrative response styles to deal with different response scenarios. To conserve space, we omitted the full prompt and detailed explanation of each component (i.e., ¡historical conversation¿, ¡current user query¿, and ¡current time¿) since they are identical to those shown in Figure 9. The explanations of these items are also shortened for the error-handling and knowledge-enhanced dialog replier.
A.2.4. Error-handling Dialog Replier
As discussed in the methodology, the error-handling pipeline aims to politely reject the user’s unmet request or further ask the user for more detailed ride-hailing intent. Specifically, we add a few instructions and provide corresponding demonstrations (i.e., how to incorporate the system error for a better response) based on the prompt of the specialized dialog replier shown in Figure 10. The added prompt content is shown in Figure 11. As shown in the prompt, the core is to add the explanation of the contradictions identified and to provide examples of contradictions to help the LLMs know how to reply in these situations.
Figure 11. The extended part of the prompt for error-handling dialog replier.
A.2.5. Knowledge-enhanced Dialog Replier
We add retrieved QA pairs into the prompt and explicitly prompt LLMs to follow these similar QA pairs to generate the response. Specifically, we achieve prompt extension (i.e., adding a task re-profile and retrieved QA pairs) based on the prompt of the specialized dialog replier shown in Figure 10. The added prompt template is shown in Figure 12. As can be seen, we provide current user query, retrieved assistant response, and corresponding similarity score in the prompt, beginning with an initial task re-profile.
Figure 12. The extended part of the prompt for knowledge-enhanced dialog replier.
A.3. APIs and Tools Lib of DiMA
This section presents all annotated APIs and tools used in DiMA, including Get_current_location, POI_search, POI_select, Route_planning and a Time tool. The following table briefly illustrates the input and output of these APIs and tools.
def Get_current_location(session_id):
    POI_list = requests.post(session_id)
    # [{displayname, lat, lng, id}, ...]
    return POI_list
def POI_search(session_id, POI_name):
    POI_list = requests.post(session_id, POI_name)
    # [{displayname, lat, lng, id}, ...]
    return POI_list
def POI_select(POI_list, Selected_POI_name):
    # {displayname, lat, lng, id}
    return Selected_POI
def Route_planning_API(session_id, SrcLat, SrcLng,
ViaLat, ViaLng, DstLat, DstLng, SrcId, ViaId, DstId):
    order_list = requests.post(session_id, ...)
    # [{dist, duration, fee, car_type}, ...]
    return Order_list
def Time_tool(date, time, current_time):
    date: "%Y-%m-%d" or "N week later, Mon" or "absent"
    time: "%H:%M:%S" or "absent"
    current_time: "%Y-%m-%d %H:%M:%S"
    return date_time # "%Y-%m-%d %H:%M:%S"
Specifically, the $session,id$ is the tag of each dialog and is also the token in the DiDi ride-hailing platform to access diverse APIs service in DiDi. When executing the function generated, the extracted spatial information (i.e., the POI name of the user’s start location, via location, or end location) and temporal information (i.e., the date and time in the user query) will be used to invoke these APIs and tools. After invocation, as discussed in Section 4.3, the corresponding ride-hailing order will be created.
A.4. Evaluation of DiMA
In this section, we will provide more details about the GPT evaluation and Human evaluation in this section, including the prompt used for GPT evaluation and how to instruct human annotator to fill the evaluation form.
A.4.1. Prompt Template of GPT Evaluation
We evaluate the performance of ride-hailing assistant from two perspective, i.e., order planning and assistant response.
Figure 13. The prompt template of GPT evaluation for order planning.
As shown in Figure 13, we explicitly prompt the LLM to return its evaluation results and corresponding reasons. Specifically, we first clearly explain each of conversation goals, e.g., confirming the order or wanting to cancel the order. Then we explain what information should be contained in the order information, e.g., start location and the price of the car type. After that, we prompt GPT-4o to evaluate if the current order information satisfies the user’s intent and to return its evaluation. Overall, in the GPT evaluation, the following spatiotemporal information and order information should be evaluated: (1) start location; (2) via location; (3) end location; (4) departure time; (5) arrival time; (6) price and car type in the order information; (7) user intent. Only if all the above information is correct will the LLM view the order creation as TRUE.
For the response generation, as shown in Figure 14, we first provide several evaluation examples in the prompt (both positive and negative samples) to cover three response scenarios. Then we provide evaluation rules, which is mainly formulated by the product manager for DiDi ride-hailing assistant, to guide LLM evaluate if current response can satisfy user’s request and align with pre-defined product rule. Specifically, the LLM will return a score (from 0 to 5), the detailed scoring criteria is as follow:
Figure 14. The prompt template of GPT evaluation for response generation.
A.4.2. Detailed Process of Human Evaluation
@@
Table 7. Study of cross-city transferability.
[…]
ROA RRA SOA SRA ROA RRA SOA SRA
Beijing $\to$ Shanghai 98.03 98.92 95.96 96.81 83.95 96.13 63.20 93.13
Shanghai $\to$ Beijing 97.73 96.36 93.26 90.03 83.36 97.21 59.09 94.32
A.5. Study of Transferability
At the initial stage, DiMA was deployed to support ride-hailing services within Beijing. To ensure its scalability to other cities, we conducted further research on cross-city transferability. As shown in Table 7, we adopted two experimental setups: (1) training on Beijing data and testing on Shanghai data, and (2) training on Shanghai data and testing on Beijing data. The experimental results demonstrate that the LLM-based ride-hailing assistant is not influenced by city-specific data, exhibiting strong scalability. Currently, DiMA is no longer restricted by city boundaries and can provide ride-hailing assistant services in any city across the country.
A.6. Case Study of DiMA
We present a case study of DiMA to illustrate how DiMA creates a ride-hailing order and provides a response given a user query. Figure 15 shows the entire dialog session.
@@ A.6. Case Study of DiMA
Figure 15. Illustrative case study. We display the concatenated prompt and the output of the spatiotemporal-aware tool-augmented order planning with the blue block. Yellow color block is used to represent the order information and the executed results. For the dialog replier selection, we use the green block to represent the output. Finally, we display response content of three dialog repliers in a orange block.

### v2

@@ 10. Conclusion
Acknowledgements.
This work was supported by the National Key R&D Program of China (Grant No.2023YFF0725004), National Natural Science Foundation of China (Grant No.92370204), the Guangzhou Basic and Applied Basic Research Program under Grant No. 2024A04J3279, Education Bureau of Guangzhou Municipality. We appreciate the guidance provided by the engineers in Didichuxing Co. Ltd.
@@ Appendix A Appendix
A.1. Simulation Environment Construction
A.1.1. Dialog Content Generation
Dialog Timestamp: 2024-7-28 19:00:00
Ride-hailing Intent: The user wants to travel to the intersection of Lou Shan Guan Road and Mao Tai Road.
Simulated Dialog Content:
<User>: I want to go to the intersection of Lou Shan Guan Road and Mao Tai Road.
<Assistant>: Okay, XiaoDi has set "the intersection of Lou Shan Guan Road and Mao Tai Road" as your drop-off point. The driver is expected to pick you up between 19:58 and 20:02 today, and it will take approximately 25 minutes to reach your destination.
<User>: Will there be traffic during this time period?
<Assistant>: Sorry, XiaoDi is unable to provide real-time traffic flow information.
A.1.2. GPT-based User Role-playing
With the generated dialog content using collected online data, we can further prompt GPT-4o to play the role of a real-world user following the dialog content. Specifically, we first provide a task profile in the prompt to guide GPT-4o in understanding the task, and then we provide the dialog content and prompt it to generate the first sentence to begin a dialog. In such a multi-dialog scenario, the response of the assistant will also be iteratively fed into the prompt to help GPT-4o understand if the ride-hailing intent has been completed. We provide detailed prompt in our GitHub project https://github.com/usail-hkust/DiMA to facilitate the understanding of GPT-based user role-playing.
A.1.3. Instruction Dataset Statistics
@@ A.2. Prompt Template of DiMA
This section presents the prompt template of every module in DiMA to facilitate a better understanding. We release the prompt template of the following five components in our GitHub project https://github.com/usail-hkust/DiMA: (1) Spatiotemporal-aware Tool-augmented Order Planning; (2) Multi-type Dialog Replier Selection; (3) Specialized Dialog Relpier; (4) Error-handling Dialog Replier; (5) Knowledge-enhanced Dialog Replier.
A.3. Evaluation of DiMA
@@
We evaluate the performance of ride-hailing assistant from two perspective, i.e., order planning and assistant response. All prompt template could be found in our GitHub project.
For order planning, we explicitly prompt the LLM to return its evaluation results and corresponding reasons. Specifically, we first clearly explain each of conversation goals, e.g., confirming the order or wanting to cancel the order. Then we explain what information should be contained in the order information, e.g., start location and the price of the car type. After that, we prompt GPT-4o to evaluate if the current order information satisfies the user’s intent and to return its evaluation. Overall, in the GPT evaluation, the following spatiotemporal information and order information should be evaluated: (1) start location; (2) via location; (3) end location; (4) departure time; (5) arrival time; (6) price and car type in the order information; (7) user intent. Only if all the above information is correct will the LLM view the order creation as TRUE.
For the response generation, we first provide several evaluation examples in the prompt (both positive and negative samples) to cover three response scenarios. Then we provide evaluation rules, which is mainly formulated by the product manager for DiDi ride-hailing assistant, to guide LLM evaluate if current response can satisfy user’s request and align with pre-defined product rule. Specifically, the LLM will return a score (from 0 to 5), the detailed scoring criteria is as follow:
A.3.1. Detailed Process of Human Evaluation
A.4. Case Study of DiMA
We present a case study of DiMA to illustrate how DiMA creates a ride-hailing order and provides a response given a user query. The figure shown in our GitHub Project https://github.com/usail-hkust/DiMA shows the entire dialog session.